### v1

# Monotonic normalized heat diffusion for distance-regular graphs with classical parameters of diameter $3$

Abstract
We prove the monotonic normalized heat diffusion property on distance-regular graphs with classical parameters of diameter $3$. Regev and Shinkar found a Cayley graph for which this property fails. On the other hand, this property has been proved on abelian Cayley graphs, graphs with $3$ distinct eigenvalues and regular bipartite graphs with $4$ distinct eigenvalues by Price, Nica and Kubo-Namba, respectively. A distance regular graph with classical parameters of diameter $3$ has $4$ distinct eigenvalues and is not necessarily bipartite or vertex transitive.
1 Introduction
In this paper, we study the heat kernel $H:[0,\infty)\times V\times V\to\mathbb{R}$ on a simple finite connected graph $G=(V,E)$. For any two vertices $u,v\in V$ and any $t\geq 0$, $H_{t}(u,v)$ can be considered as the probability that a continuous-time random walker starting at vertex $u$, is at vertex $v$ at time $t$. For any two distinct vertices $u,v\in V$, we consider the function
$t\mapsto r_{t}(u,v):=\frac{H_{t}(u,v)}{H_{t}(u,u)}.$
Clearly, we have
$r_{0}(u,v)=0,\,\,\text{and}\,\,\lim_{t\to\infty}r_{t}(u,v)=1.$ (1)
It is natural to ask whether the above function is monotonically non-decreasing or not. This question has been considered by Regev and Shinkar [12]. We will refer to this property as the monotonic normalized heat diffusion property, or the MNHD property for short, following [8, 10].
Due to (1), a necessary condition for the MNHD property is that $r_{t}(u,v)\leq 1$ holds for any $t\geq 0$. However, there do exist graphs for which the function $r_{t}(u,v)$ exceeds 1 for certain vertices $u,v$ at certain time $t$. This is observed by Regev and Shinkar [12, Proposition A.1]. Regev and Shinkar mentioned that this observation was brought to their attention by Cheeger. They also provided an explicit example which is a $4$-regular graph with $10$ vertices [12, Figure 2].
A sufficient condition to ensure that $r_{t}(u,v)\leq 1$ is a heat homogeneity, that is, the heat kernel has constant diagonal, which means that the value of $H_{t}(u,u)$ is independent of the choice of $u$ (see [12, Section 1] and [10, Section 1]). In fact, the heat homogeneity holds if and only if the graph is walk-regular [10, Theorem A.1]. Recall that a graph is walk-regular if for every $\ell\geq 2$, the number of closed walks of length $\ell$ starting from and ending at every vertex is constant. Typical walk-regular graphs includes vertex-transitive graphs, distance-regular graphs, and regular graphs with at most $4$ distinct eigenvalues.
Peres asked in 2013 (see [12, Section 1]) whether the MNHD property holds for all vertex transitive graphs or not. Regev and Shinkar [12] showed that there exists a Cayley graph for which the MNHD property fails. This provides a negative answer to Peres’ question. On the other hand, Price [11] proved that the MNHD property does hold for all finite abelian Cayley graphs.
Nica [10] showed that the MNHD property holds for graphs with $3$ distinct Laplacian eigenvalues. We stress that the result of Nica does not require the graph to be regular. Kubo and Namba [8] confirmed the MNHD property for regular bipartite graphs with $4$ distinct eigenvalues. It is natural to ask whether the MNHD property holds for all regular graphs with at most $4$ distinct eigenvalues.
Distance-regular graphs are introduced by Biggs [1] as a natural extension of distance-transitive graphs. A graph is distance-regular if for any two vertices $u,v$ at distance $h$, the number of vertices that are at distance $i$ to $u$ and at distance $j$ to $v$ only depends on the three numbers $h,i$ and $j$. Every distance-regular grpah has diameter $D$ if and only if it has $D+1$ distinct eigenvalues.
For a regular graph $G$ with $4$ distinct eigenvalue, van Dam and Haemers [5] showed that $G$ is distance-regular if and only if the number of vertices at distance $2$ from each given vertex satisfies an expression in terms of eigenvalues of $G$.
We identify a class of distance-regular graphs with $4$ distinct eigenvalues that are not necessarily bipartite, for which the MNHD property holds true. Our main result states as follows.
Theorem 1.
Let $G=(V,E)$ be a distance-regular graph with classical parameters of diameter $3$. Then $G$ has the monotonic normalized heat diffusion (MNHD) property.
Most distance-regular graphs related to classical groups and groups of Lie type have classical parameters. Lots of interesting graphs constructed from algebraic structures are either distance-regular graphs with classical parameters (see, e.g., [7] and [3, Tables 6.1 and 6.2]) or are closely related to distance-regular graphs with classical parameters (see, e.g., [4] and [6, Section 3.2]). We refer to Subsection 2.3 for the precise definition of a distance-regular graph with classical parameters. Although there exist various classification results for distance-regular graphs with classical parameters of large diameter (see, e.g. [6]), relatively fewer results are known for the diameter $3$ case. We believe that our investigation into the monotonic normalized heat diffusion (MNHD) property sheds new light on the structure of distance-regular graphs with classical parameters.
In [2], Biggs classified the distance-regular graphs of diameter 3 into three categories: antipodal, bipartite and primitive. We show in Section 5 that the MNHD property holds for all antipodal distance-regular graphs of diameter $3$. Notice that the bipartite case follows from the result of Kubo and Namba [8].
2 Preliminaries
In this section, we collect basics about graph Laplacian, heat kernel, and distance-regular graphs with classical parameters.
2.1 Graph Laplacian and heat kernels
Let $d\geq 1$ be an integer and $G=(V,E)$ a simple finite connected graph with vertex set $V$ and edge set $E$. Suppose that $|V|=n\geq 2$. The graph Laplacian $L$ is a symmetric $n\times n$ matrix $L=(L(x,y))_{x,y\in V}$ defined by $L=\mathrm{Deg}-A$, where $\mathrm{Deg}$ is the diagonal vertex degree matrix and $A$ is the adjacency matrix. The constant vector $\mathbf{1}=(1,1,...,1)^{T}$ is an eigenvector of $L$ corresponding to the eigenvalue 00. Since $G$ is connected, the eigenvalue 00 is simple by the Perron-Frobenius theorem. Let us denote by $\mathrm{sp}(L)$ the set of all distinct eigenvalues of $L$. The graph Laplacian has the following spectral decomposition
$L=\sum_{\lambda\in\mathrm{sp}(L)}\lambda P_{\lambda},$ (2)
[…]
The heat kernel $H_{t}$ on the graph $G$ is defined in matrix form as $H_{t}:=e^{-tL}$. Accordingly, $H_{t}$ has the following spectral decomposition
$H_{t}=\sum_{\lambda\in\mathrm{sp}(L)}e^{-t\lambda}P_{\lambda}\ ,$ (3)
for all $t\geq 0$. For any two distinct vertices $u,v\in V$, we define the function $h_{u,v}:[0,\infty)\to\mathbb{R}$ as follows,
$h_{u,v}(t):=H_{t}^{\prime}(u,v)H_{t}(u,u)-H_{t}(u,v)H_{t}^{\prime}(u,u).$ (4)
Here $H_{t}^{\prime}$ stands for the derivative of $H_{t}$ with respect to $t$. The MNHD property is equivalent to that $h_{u,v}(t)\geq 0$ for any $t\geq 0$. Nica observed that this is always true at $t=0$.
Lemma 1 ([10, Lemma 3.2]).
For any two distinct vertices $u,v$, it holds that $h_{u,v}(0)\geq 0$.
Let us list all $\lambda\in\mathrm{sp}(L)$ as follows:
$0<\lambda_{1}<\lambda_{2}<\cdots<\lambda_{s}.$ (5)
Then, we have the following expression of $h_{u,v}(t)$, see, e.g., [8, (3.2)].
Lemma 2.
For any two distinct vertices $u,v$ and $t\geq 0$, we have
$h_{u,v}(t)=\sum_{i=1}^{s}\frac{\lambda_{i}}{n}e^{-t\lambda_{i}}\Delta_{i}(u,v)+\sum_{1\leq i<j\leq s}(\lambda_{j}-\lambda_{i})e^{-t(\lambda_{i}+\lambda_{j})}\Delta_{ij}(u,v),$ (6)
where for any $i,j=1,\ldots,s$,
$\Delta_{i}(u,v)$ $:=P_{\lambda_{i}}(u,u)-P_{\lambda_{i}}(u,v),$
$\Delta_{ij}(u,v)$ $:=P_{\lambda_{i}}(u,v)P_{\lambda_{j}}(u,u)-P_{\lambda_{j}}(u,v)P_{\lambda_{i}}(u,u).$
Proof.
This follows from (3) by a straightforward calculation. ∎
2.2 Heat kernels on regular graphs with four distinct eigenvalues
We consider a simple finite connected $d$-regular graph $G=(V,E)$ with $4$ distinct Laplacian eigenvalues, i.e., $s=3$ in (5). In this particular setting, we have the following explicit expressions for the functions $\Delta_{i}(u,v)$ and $\Delta_{ij}(u,v)$ due to Kubo and Namba [8].
Lemma 3 ([8, Lemmas 3 and 4]).
Let $G=(V,E)$ be a simple finite connected $d$ regular graph with $4$ distinct Laplacian eigenvalues $0<\lambda_{1}<\lambda_{2}<\lambda_{3}.$ Denote by
$C_{1}=\frac{1}{(\lambda_{2}-\lambda_{1})(\lambda_{3}-\lambda_{1})},\,\,C_{2}=\frac{1}{(\lambda_{3}-\lambda_{2})(\lambda_{1}-\lambda_{2})},\,\,\text{and}\,\,\,C_{3}=\frac{1}{(\lambda_{1}-\lambda_{3})(\lambda_{2}-\lambda_{3})}.$
Then, for any two distinct vertices $u,v\in V$, we have
$\Delta_{1}(u,v)$ $=C_{1}\{d^{2}+d-L^{2}(u,v)+(L(u,v)-d)(\lambda_{2}+\lambda_{3})+\lambda_{2}\lambda_{3}\},$ (7)
$\Delta_{2}(u,v)$ $=C_{2}\{d^{2}+d-L^{2}(u,v)+(L(u,v)-d)(\lambda_{1}+\lambda_{3})+\lambda_{1}\lambda_{3}\},$ (8)
$\Delta_{3}(u,v)$ $=C_{3}\{d^{2}+d-L^{2}(u,v)+(L(u,v)-d)(\lambda_{1}+\lambda_{2})+\lambda_{1}\lambda_{2}\},$ (9)
and for $\{i,j,k\}=\{1,2,3\}$,
$\Delta_{ij}(u,v)=C_{i}C_{j}(\lambda_{j}-\lambda_{i})$$\bigg{\{}\left(d+\frac{1-n}{n}\lambda_{k}\right)L^{2}(u,v)$
$-\left(d^{2}+d+\frac{1-n}{n}\lambda_{k}^{2}\right)L(u,v)-\frac{1}{n}\left((d^{2}+d)\lambda_{k}-d\lambda_{k}^{2}\right)\bigg{\}}.$ (10)
Proof.
The expressions (7)-(9) are obtained in [8, Proof of Lemma 3] and equation (10) is a rearrangement of [8, (3.4)]. For readers’ convenience, we briefly recall the key steps of the proof. Solving the equations $I=P_{0}+P_{\lambda_{1}}+P_{\lambda_{2}}+P_{\lambda_{3}}$ and $L^{k}=\lambda_{1}^{k}P_{\lambda_{1}}+\lambda_{2}^{k}P_{\lambda_{2}}+\lambda_{3}^{k}P_{\lambda_{3}}$, $k=1,2$ leads to the following expressions for the projection operators:
$P_{\lambda_{1}}$ $=C_{1}\big{\{}L^{2}-(\lambda_{2}+\lambda_{3})L+\lambda_{2}\lambda_{3}(I-P_{0})\big{\}},$
$P_{\lambda_{2}}$ $=C_{2}\big{\{}L^{2}-(\lambda_{1}+\lambda_{3})L+\lambda_{1}\lambda_{3}(I-P_{0})\big{\}},$
$P_{\lambda_{3}}$ $=C_{3}\big{\{}L^{2}-(\lambda_{1}+\lambda_{2})L+\lambda_{1}\lambda_{2}(I-P_{0})\big{\}}.$
Then Lemma 3 follows by a straightforward calculation, noticing that $L(u,u)=d$ and $L^{2}(u,u)=d^{2}+d$. ∎
2.3 Distance-regular graph with classical parameters
A connected graph $G=(V,E)$ of diameter $D$ is distance-regular if and only if it is regular of degree $d$, and if there exist constants $c_{i},a_{i},b_{i}$, known as intersection numbers, such that for all $i=0,1,\dots,D$, and for any two vertices $x,y\in V$ at distance $i=d(x,y)$, among the neighbors of $y$, there are $c_{i}$ at distance $i-1$ from $x$, $a_{i}$ at distance $i$, and $b_{i}$ at distance $i+1$. It follows that $d=b_{0}$, and $a_{i}=d-b_{i}-c_{i}$ for all $i=0,1,\dots,D$. These constants are commonly listed in the so-called intersection array
$\{b_{0},b_{1},\dots,b_{D-1};c_{1},c_{2},\dots,c_{D}\}.$
Observe that $b_{D}=0$ and $c_{0}=0$ are excluded from this array, while $c_{1}=1$ is included, and all numbers in the intersection array are positive integers. Furthermore, the number of vertices can be derived from the intersection array. Indeed, by counting edges $yz$ with $d(x,y)=i$ and $d(x,z)=i+1$ for each vertex $x$, it has a constant number of vertices $n_{i}$ at a given distance $i$. Specifically, it follows from $n_{0}=1$ and $n_{i+1}=\frac{b_{i}n_{i}}{c_{i+1}}$ for all $i=0,1,\dots,D-1$ that the number of vertices $n=n_{0}+n_{1}+\cdots+n_{D}$.
The distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$ are the distance-regular graphs whose intersection numbers can be expressed by the four parameters: diameter $D$ and numbers $b$, $\alpha$, $\beta$, in the following way,
$b_{i}$ $=\left(\begin{bmatrix}D\\ 1\\ \end{bmatrix}-\begin{bmatrix}i\\ 1\\ \end{bmatrix}\right)\left(\beta-\alpha\begin{bmatrix}i\\ 1\\ \end{bmatrix}\right)\;\;,i=0,1,\dots,D-1,$ (11)
$c_{i}$ $=\begin{bmatrix}i\\ 1\\ \end{bmatrix}\left(1+\alpha\begin{bmatrix}i-1\\ 1\\ \end{bmatrix}\right)\;\;,i=1,2,\dots,D,$ (12)
where $\begin{bmatrix}j\\ 1\\ \end{bmatrix}=1+b+b^{2}+\cdots+b^{j-1}$ is the Gaussian binomial coefficient with basis $b$. To make $b_{i},c_{i}$ feasible parameters, there are necessary restrictions on the numbers $b,\alpha,\beta$. For example, the number $b$ must be an integer not equal to 00 or $-1$ [3, Proposition 6.2.1]. See also Lemma 4 below.
Proposition 1 ([3, Corollary 8.4.2]).
Let $G$ be a distance-regular graph with classical parameters $(D,b,\alpha,\beta)$. Then the eigenvalues of its adjacent matrix are
$\theta_{i}=\frac{b_{i}}{b^{i}}-\begin{bmatrix}i\\ 1\\ \end{bmatrix},\ \ i=0,1,\cdots,D,$
where $b_{i},i=0,1,\ldots,D$ are given by (11).
For more discussion on distance-regular graphs, we refer to [3] and [6].
2.4 Distance-regular graphs with classical parameters of diameter 3
In this subsection, we collect basics of distance-regular graphs with classical parameters $(D,b,\alpha,\beta)$ of diameter $D=3$.
First, we derive from (11) and (12) the intersection numbers as follows.
$\begin{array}[]{ll}b_{0}=d=(1+b+b^{2})\beta,&b_{1}=(b+b^{2})(\beta-\alpha),\\ b_{2}=b^{2}(\beta-\alpha(1+b)),&b_{3}=0,\\ c_{0}=0,&c_{1}=1,\\ c_{2}=(1+b)(1+\alpha),&c_{3}=(1+b+b^{2})(1+\alpha(1+b)),\\ a_{0}=0,&a_{1}=-1+b\alpha+b^{2}\alpha+\beta,\\ a_{2}=(1+b)(-1-\alpha+b^{2}\alpha+\beta),&a_{3}=(1+b+b^{2})(-1-\alpha-b\alpha+\beta).\\ \end{array}$ (13)
Then, we calculate the vertex number $n$ as below,
$n=$ $1+d+\frac{db_{1}}{c_{2}}+\frac{db_{1}b_{2}}{c_{2}c_{3}}$
$=$ $1+d+\frac{db(\beta-\alpha)}{1+\alpha}+\frac{db(\beta-\alpha)b_{2}}{(1+\alpha)c_{3}}.$ (14)
Applying Proposition 1, we obtain the eigenvalues of its adjacent matrix as follows,
$\theta_{0}=d,\theta_{1}=(1+b)(\beta-\alpha)-1,\theta_{2}=\beta-(1+b)(1+\alpha),\,\text{and}\,\,\theta_{3}=-(1+b+b^{2}),$ (15)
and, accordingly, the non-zero Laplacian eigenvalues $\gamma_{i}=d-\theta_{i}$ listed below,
$\gamma_{1}=1+\alpha+b\alpha+b^{2}\beta,\gamma_{2}=(1+b)(1+\alpha+b\beta),\,\text{and}\,\,\gamma_{3}=(1+b+b^{2})(1+\beta).$ (16)
In the following, we derive some necessary constraints from (13) on the parameters $b,\alpha,\beta$ in case $D=3$.
Lemma 4.
Let $G=(V,E)$ be a distance-regular graph with classical parameters $(3,b,\alpha,\beta)$. Then we have the following constraints:
  • (i)
    If $b\geq 1$, then $\alpha\geq 0$.
  • (ii)
    If $b\leq-2$, then $\alpha<-1$.
  • (iii)
    $\beta\geq 1+(1+b)\alpha\geq 1$.
Proof.
Since $c_{2}=(1+b)(1+\alpha)>0$, we obtain $\alpha<-1$ if $b\leq-2$. Moreover, since $b$ is an integer, we read from $c_{2}$ that $(1+b)\alpha$ is an integer. Via $c_{3}=(1+b+b^{2})(1+\alpha(1+b))>0$, we derive that $\alpha(1+b)>-1$ if $b\geq 1$. Therefore, we have $\alpha(1+b)\geq 0$ if $b\geq 1$. This proves $(i)$. The constraint $\beta\geq 1+(1+b)\alpha$ follows from the facts $a_{3}=(1+b+b^{2})(-1-\alpha-b\alpha+\beta)\geq 0$. Applying $(i)$ and $(ii)$, we further deduce $\beta\geq 1+(1+b)\alpha\geq 1$. This completes the proof. ∎
Remark 1.
Lemma 4 $(i)$ is taken from Lv and Koolen [9]. In fact, Lemma 4 $(i)$ and $(ii)$ hold for any distance-regular graph with classical parameters of diameter $D\geq 3$. The proof of Lemma 4 $(i)$ and $(ii)$ relies only on the expressions of $c_{2}$ and $c_{3}$ in (13), which holds for any diameter $D\geq 3$.
3 The MNHD property for regular graphs with four distinct eigenvalues
In this section, we derive a sufficient condition for the MNHD property to hold on regular graphs with four distinct eigenvalues.
Our main Theorem 2 in this section relies on the following key observation.
Lemma 5.
For any three distinct numbers $\lambda_{i},\lambda_{j},\lambda_{k}\in\mathbb{R}$ with $\{i,j,k\}=\{1,2,3\}$ and any number $L(u,v),L^{2}(u,v)\in\mathbb{R}$, let $\Delta_{i}(u,v),i=1,2,3$ and $\Delta_{ij}(u,v)$ be defined in terms of $\lambda_{i},\lambda_{j},\lambda_{k},L(u,v),L^{2}(u,v)$ as in (7)-(10). Then it holds that
$\frac{\lambda_{i}\lambda_{j}}{n}\Delta_{i}(u,v)+\frac{\lambda_{j}\lambda_{i}}{n}\Delta_{j}(u,v)+\frac{\lambda_{k}(\lambda_{i}+\lambda_{j}-\lambda_{k})}{n}\Delta_{k}(u,v)$
$\hskip 170.71652pt-(\lambda_{k}-\lambda_{i})(\lambda_{k}-\lambda_{j})(\Delta_{ik}(u,v)+\Delta_{jk}(u,v))$
$=$ $L^{2}(u,v)-L(u,v)(\lambda_{i}+\lambda_{j}).$ (17)
Remark 2.
Notice that in identity (17), the three distinct numbers $\lambda_{i},\lambda_{j},\lambda_{k}\in\mathbb{R}$ are not required to be eigenvalues of the graph Laplacian $L$; the two numbers $L(u,v)$ and $L^{2}(u,v)$ are not required to be obtained from the graph Laplacian $L$ and its square $L^{2}$. We only use the formal definitions of $\Delta_{i}$ and $\Delta_{ij}$ in (7)-(10).
Proof.
We break down the calculations into two parts. First, we consider the terms involving $\Delta_{i}(u,v)$, $\Delta_{j}(u,v)$, and $\Delta_{k}(u,v)$. We calculate by inserting (7)-(9) that
$\frac{\lambda_{i}\lambda_{j}}{n}\Delta_{i}(u,v)+\frac{\lambda_{j}\lambda_{i}}{n}\Delta_{j}(u,v)+\frac{\lambda_{k}(\lambda_{i}+\lambda_{j}-\lambda_{k})}{n}\Delta_{k}(u,v)$
$=$ $\frac{\lambda_{i}\lambda_{j}\left(d+d^{2}-L^{2}(u,v)+(L(u,v)-d)(\lambda_{j}+\lambda_{k})+\lambda_{j}\lambda_{k}\right)}{n(\lambda_{i}-\lambda_{j})(\lambda_{i}-\lambda_{k})}$
$+\frac{\lambda_{i}\lambda_{j}\left(d+d^{2}-L^{2}(u,v)+(L(u,v)-d)(\lambda_{i}+\lambda_{k})+\lambda_{i}\lambda_{k}\right)}{n(\lambda_{j}-\lambda_{i})(\lambda_{j}-\lambda_{k})}$
$+\frac{\lambda_{k}(\lambda_{i}+\lambda_{j}-\lambda_{k})\left(d+d^{2}-L^{2}(u,v)+(L(u,v)-d)(\lambda_{i}+\lambda_{j})+\lambda_{i}\lambda_{j}\right)}{n(\lambda_{k}-\lambda_{i})(\lambda_{k}-\lambda_{j})}$
$=$ $\frac{(d+d^{2}-L^{2}(u,v))\mathrm{I}+(L(u,v)-d)\mathrm{II}+\mathrm{III}}{n(\lambda_{i}-\lambda_{j})(\lambda_{i}-\lambda_{k})(\lambda_{j}-\lambda_{k})},$
where we have
$\mathrm{I}:=$ $\lambda_{i}\lambda_{j}(\lambda_{j}-\lambda_{k})-\lambda_{i}\lambda_{j}(\lambda_{i}-\lambda_{k})+\lambda_{k}(\lambda_{i}+\lambda_{j}-\lambda_{k})(\lambda_{i}-\lambda_{j})$
$=$ $-(\lambda_{i}-\lambda_{j})(\lambda_{i}-\lambda_{k})(\lambda_{j}-\lambda_{k}),$
$\mathrm{II}:=$ $\lambda_{i}\lambda_{j}(\lambda_{j}^{2}-\lambda_{k}^{2})-\lambda_{i}\lambda_{j}(\lambda_{i}^{2}-\lambda_{k}^{2})+\lambda_{k}(\lambda_{i}+\lambda_{j}-\lambda_{k})(\lambda_{i}^{2}-\lambda_{j}^{2})$
$=$ $-(\lambda_{i}-\lambda_{j})(\lambda_{i}-\lambda_{k})(\lambda_{j}-\lambda_{k})(\lambda_{i}+\lambda_{j}),$
$\mathrm{III}:=$ $\lambda_{i}\lambda_{j}^{2}\lambda_{k}(\lambda_{j}-\lambda_{k})-\lambda_{i}^{2}\lambda_{j}\lambda_{k}(\lambda_{i}-\lambda_{k})+\lambda_{i}\lambda_{j}\lambda_{k}(\lambda_{i}+\lambda_{j}-\lambda_{k})(\lambda_{i}-\lambda_{j})$
$=$ $0$
Putting together yields that
$\frac{\lambda_{i}\lambda_{j}}{n}\Delta_{i}(u,v)+\frac{\lambda_{j}\lambda_{i}}{n}\Delta_{j}(u,v)+\frac{\lambda_{k}(\lambda_{i}+\lambda_{j}-\lambda_{k})}{n}\Delta_{k}(u,v)$
$=$ $-\frac{1}{n}\left\{d+d^{2}-L^{2}(u,v)+(L(u,v)-d)(\lambda_{i}+\lambda_{j})\right\}.$ (18)
Next, we consider the terms involving $\Delta_{ik}(u,v)$ and $\Delta_{jk}(u,v)$. By inserting (10), we calculate that
$-(\lambda_{k}-\lambda_{i})(\lambda_{k}-\lambda_{j})(\Delta_{ik}(u,v)+\Delta_{jk}(u,v))$
$=$ $\frac{(nd+(1-n)\lambda_{j})L^{2}(u,v)-(n(d^{2}+d)+(1-n)\lambda_{j}^{2})L(u,v)-(d^{2}+d)\lambda_{j}+d\lambda_{j}^{2}}{n(\lambda_{i}-\lambda_{j})}$
$+\frac{(nd+(1-n)\lambda_{i})L^{2}(u,v)-(n(d^{2}+d)+(1-n)\lambda_{i}^{2})L(u,v)-(d^{2}+d)\lambda_{i}+d\lambda_{i}^{2}}{n(\lambda_{j}-\lambda_{i})}$
$=$ $\frac{1}{n}\left\{-(1-n)\left(L^{2}(u,v)-L(u,v)(\lambda_{i}+\lambda_{j})\right)+d^{2}+d-d(\lambda_{i}+\lambda_{j})\right\}.$ (19)
Combining (18) and (19) leads to the identity (5). ∎
Next, we establish a sufficient condition for a regular graph with four distinct eigenvalues to have MNHD property.
Theorem 2.
Let $G$ be a regular graph with four distinct Laplacian eigenvalues
$0<\lambda_{1}<\lambda_{2}<\lambda_{3}.$
Let $u,v$ be two distinct vertices. Suppose that $\Delta_{i}(u,v)\geq 0,i=1,2,3$, and one of the following conditions holds,
  • (i)
    $L^{2}(u,v)-L(u,v)(\lambda_{2}+\lambda_{3})\geq 0$, and $\Delta_{12}(u,v)\geq 0$,
  • (ii)
    $L^{2}(u,v)-L(u,v)(\lambda_{1}+\lambda_{2})\geq 0$, $\Delta_{13}(u,v)\leq 0$, and $\lambda_{1}+\lambda_{2}-\lambda_{3}\geq 0$,
  • (iii)
    $L^{2}(u,v)-L(u,v)(\lambda_{1}+\lambda_{3})\geq 0$, $\Delta_{23}(u,v)\geq 0$, and $\lambda_{1}+\lambda_{2}-\lambda_{3}\geq 0$.
Then the function $r_{t}(u,v):=\frac{H_{t}(u,v)}{H_{t}(u,u)}$ is monotonically non-decreasing in $t$.
Proof.
Recall that the monotonic property of $r_{t}(u,v)$ in $t$ is equivalent to the non-negativity of the function $h_{u,v}(t)$ defined in (4) and (6).
Case $(i)$: Consider the function $F(t):=e^{t(\lambda_{2}+\lambda_{3})}h_{u,v}(t)$. Employing Lemma 2, we derive
$F(t)=$ $\frac{\lambda_{1}}{n}\Delta_{1}(u,v)e^{t(\lambda_{2}+\lambda_{3}-\lambda_{1})}+\frac{\lambda_{2}}{n}\Delta_{2}(u,v)e^{t\lambda_{3}}+\frac{\lambda_{3}}{n}\Delta_{3}(u,v)e^{t\lambda_{2}}$
$+(\lambda_{2}-\lambda_{1})\Delta_{12}(u,v)e^{t(\lambda_{3}-\lambda_{1})}+(\lambda_{3}-\lambda_{1})\Delta_{13}(u,v)e^{t(\lambda_{2}-\lambda_{1})}$
$+(\lambda_{3}-\lambda_{2})\Delta_{23}(u,v).$
Then, we calculate
$F^{\prime}(t)=$ $\frac{\lambda_{1}}{n}\Delta_{1}(u,v)(\lambda_{2}+\lambda_{3}-\lambda_{1})e^{t(\lambda_{2}+\lambda_{3}-\lambda_{1})}+\frac{\lambda_{2}\lambda_{3}}{n}\Delta_{2}(u,v)e^{t\lambda_{3}}$
$+\frac{\lambda_{3}\lambda_{2}}{n}\Delta_{3}(u,v)e^{t\lambda_{2}}+(\lambda_{3}-\lambda_{1})(\lambda_{2}-\lambda_{1})\Delta_{12}(u,v)e^{t(\lambda_{3}-\lambda_{1})}$
$+(\lambda_{2}-\lambda_{1})(\lambda_{3}-\lambda_{1})\Delta_{13}(u,v)e^{t(\lambda_{2}-\lambda_{1})}.$
Extracting the factor $e^{t(\lambda_{2}-\lambda_{1})}$, we derive
$F^{\prime}(t)=$ $e^{t(\lambda_{2}-\lambda_{1})}\bigg{(}\frac{\lambda_{1}}{n}\Delta_{1}(u,v)(\lambda_{2}+\lambda_{3}-\lambda_{1})e^{t\lambda_{3}}+\frac{\lambda_{3}\lambda_{2}}{n}\Delta_{2}(u,v)e^{t(\lambda_{3}-\lambda_{2}+\lambda_{1})}$
$+\frac{\lambda_{3}\lambda_{2}}{n}\Delta_{3}(u,v)e^{t\lambda_{1}}+(\lambda_{3}-\lambda_{1})(\lambda_{2}-\lambda_{1})\Delta_{12}(u,v)e^{t(\lambda_{3}-\lambda_{2})}$
$+(\lambda_{3}-\lambda_{1})(\lambda_{2}-\lambda_{1})\Delta_{13}(u,v)\bigg{)}.$
Since $\Delta_{i}(u,v)\geq 0,i=1,2,3$ and $\Delta_{12}(u,v)\geq 0$, we estimate
$e^{-t(\lambda_{2}-\lambda_{1})}F^{\prime}(t)\geq$ $\frac{\lambda_{1}(\lambda_{2}+\lambda_{3}-\lambda_{1})}{n}\Delta_{1}(u,v)+\frac{\lambda_{3}\lambda_{2}}{n}\Delta_{2}(u,v)+\frac{\lambda_{3}\lambda_{2}}{n}\Delta_{3}(u,v)$
$+(\lambda_{3}-\lambda_{1})(\lambda_{2}-\lambda_{1})(\Delta_{12}(u,v)+\Delta_{13}(u,v))$
$=$ $L^{2}(u,v)-L(u,v)(\lambda_{2}+\lambda_{3}).$
We have used the fact that $\Delta_{ij}(u,v)=-\Delta_{ji}(u,v)$ and Lemma 5 in the above estimate. By assumption that $L^{2}(u,v)-L(u,v)(\lambda_{2}+\lambda_{3})\geq 0$, we derive
$F(t)\geq F(0)=h_{u,v}(0)\geq 0,\ \text{for any}\ t\geq 0,$
where we used Lemma 1. Therefore, we obtain $h_{u,v}(t)\geq 0,\text{for all}\ t\geq 0.$
Case $(ii)$: We consider the function $G(t)=e^{t(\lambda_{1}+\lambda_{2})}h_{u,v}(t)$. Employing Lemma 2, we derive
$G(t)=$ $\frac{\lambda_{1}}{n}\Delta_{1}(u,v)e^{t\lambda_{2}}+\frac{\lambda_{2}}{n}\Delta_{2}(u,v)e^{t\lambda_{1}}+\frac{\lambda_{3}}{n}\Delta_{3}(u,v)e^{t(\lambda_{1}+\lambda_{2}-\lambda_{3})}$
$+(\lambda_{2}-\lambda_{1})\Delta_{12}(u,v)+(\lambda_{3}-\lambda_{1})\Delta_{13}(u,v)e^{t(\lambda_{2}-\lambda_{3})}$
$+(\lambda_{3}-\lambda_{2})\Delta_{23}(u,v)e^{t(\lambda_{1}-\lambda_{3})}.$
Similarly as in Case $(i)$, we estimate
$e^{-t(\lambda_{1}-\lambda_{3})}G^{\prime}(t)=\geq$ $\frac{\lambda_{1}\lambda_{2}}{n}\Delta_{1}(u,v)+\frac{\lambda_{2}\lambda_{1}}{n}\Delta_{2}(u,v)+\frac{\lambda_{3}(\lambda_{1}+\lambda_{2}-\lambda_{3})}{n}\Delta_{3}(u,v)$
$+(\lambda_{3}-\lambda_{1})(\lambda_{2}-\lambda_{3})(\Delta_{13}(u,v)+\Delta_{23}(u,v))$
$=$ $L^{2}(u,v)-L(u,v)(\lambda_{1}+\lambda_{2}).$
The last identity is derived from Lemma 5. In particular, we mention that our assumption $\lambda_{1}+\lambda_{2}-\lambda_{3}\geq 0$ is used here.
Therefore, we deduce that
$G(t)\geq G(0)=h_{u,v}(0)\geq 0,\ \text{for any}\ t\geq 0.$
Hence, we have $h_{u,v}(t)\geq 0$, for any $t\geq 0$.
Case $(iii)$: We consider the function $H(t)=e^{t(\lambda_{1}+\lambda_{3})}h_{u,v}(t)$. Employing Lemma 2, we derive
$H(t)=$ $\frac{\lambda_{1}}{n}\Delta_{1}(u,v)e^{t\lambda_{3}}+\frac{\lambda_{2}}{n}\Delta_{2}(u,v)e^{t(\lambda_{1}+\lambda_{3}-\lambda_{2})}+\frac{\lambda_{3}}{n}\Delta_{3}(u,v)e^{t\lambda_{1}}$
$+(\lambda_{2}-\lambda_{1})\Delta_{12}(u,v)e^{t(\lambda_{3}-\lambda_{2})}+(\lambda_{3}-\lambda_{1})\Delta_{13}(u,v)$
$+(\lambda_{3}-\lambda_{2})\Delta_{23}(u,v)e^{t(\lambda_{1}-\lambda_{2})}.$
Similarly as in Case $(ii)$, we estimate
$e^{-t(\lambda_{3}-\lambda_{2})}H^{\prime}(t)$
$=$ $\frac{\lambda_{1}\lambda_{3}}{n}\Delta_{1}(u,v)e^{t\lambda_{2}}+\frac{\lambda_{2}(\lambda_{1}+\lambda_{3}-\lambda_{2})}{n}\Delta_{2}(u,v)e^{t\lambda_{1}}+\frac{\lambda_{3}\lambda_{1}}{n}\Delta_{3}(u,v)e^{t(\lambda_{1}+\lambda_{2}-\lambda_{3})}$
$+(\lambda_{3}-\lambda_{2})(\lambda_{1}-\lambda_{2})\Delta_{23}(u,v)e^{t(\lambda_{1}-\lambda_{3})}+\left.(\lambda_{2}-\lambda_{1})(\lambda_{3}-\lambda_{2})\Delta_{12}(u,v)\right)$
$\geq$ $\frac{\lambda_{1}\lambda_{3}}{n}\Delta_{1}(u,v)+\frac{\lambda_{2}(\lambda_{1}+\lambda_{3}-\lambda_{2})}{n}\Delta_{2}(u,v)+\frac{\lambda_{3}\lambda_{1}}{n}\Delta_{3}(u,v)$
$+(\lambda_{3}-\lambda_{2})(\lambda_{1}-\lambda_{2})\Delta_{23}(u,v)+(\lambda_{2}-\lambda_{1})(\lambda_{3}-\lambda_{2})\Delta_{12}(u,v)$
$=$ $L^{2}(u,v)-L(u,v)(\lambda_{1}+\lambda_{3}).$
Notice that we have used $\lambda_{1}+\lambda_{2}-\lambda_{3}\geq 0$ in the inequality above and Lemma 5 in the equality. It follows that
$H(t)\geq H(0)=h_{u,v}(0)\geq 0,\ \text{for any}\ t\geq 0.$
Hence, we have $h_{u,v}(t)\geq 0$, for any $t\geq 0$. This completes the proof. ∎
Remark 3.
Observe from $L=dI-A$ that
$L^{2}(x,y)$ $=d^{2}I(x,y)-2dA(x,y)+A^{2}(x,y)$
=$=\begin{cases}d^{2}+d,&\text{if $x=y$,}\\ -2dA(x,y)+\sum_{w\neq x,y}A(x,w)A(w,y),&\text{if $x\neq y$.}\end{cases}$ (20)
Therefore, for any non-adjacent vertices $u,v$, the condition $L^{2}(u,v)-L(u,v)(\lambda_{i}+\lambda_{j})\geq 0$ in Theorem 2 always holds.
Remark 4.
Let $G$ be a regular bipartite graph with four distinct eigenvalues $0<\lambda_{1}<\lambda_{2}<\lambda_{3}$. Then the conditions $L^{2}(u,v)-L(u,v)(\lambda_{i}+\lambda_{j})\geq 0$ and $\lambda_{1}+\lambda_{2}-\lambda_{3}\geq 0$ in Theorem 2 all holds true. Indeed, we have $L^{2}(u,v)=-2d$ for adjacent vertices $u$ and $v$ from (3), and
$\lambda_{1}=d-\sqrt{d-\lambda},\lambda_{2}=d+\sqrt{d-\lambda},\lambda_{3}=2d,$
where $\lambda:=2d(d-1)/(n-2)$ with $n$ being the number of vertices, see [8, Propostions 3 and 4]. Consequently, we obtain $\lambda_{i}+\lambda_{j}\geq 2d$ for any $\{i,j\}\in\{1,2,3\}$ and $\lambda_{1}+\lambda_{2}=\lambda_{3}$.
4 Proof of Theorem 1
In this section, we give the proof of our main Theorem 1. We first figure out the entries of matrix $L^{2}$ for a distance-regular graph.
Proposition 2.
Let $G=(V,E)$ be a distance-regular graph. Then we have for any $x,y\in V$ that
$L^{2}\left(x,y\right)=\begin{cases}d^{2}+d,&\text{if}\;d\left(x,y\right)=0;\\ -2d+a_{1},&\text{if}\;d\left(x,y\right)=1;\\ c_{2},&\text{if}\;d\left(x,y\right)=2;\\ 0,&\text{if}\;d\left(x,y\right)\geq 3.\\ \end{cases}$ (21)
Proof.
This follows directly from (3). We only need to observe that
$\sum_{w\neq x,y}A(x,w)A(w,y)=\begin{cases}a_{1},&\text{if $d(x,y)=1$};\\ c_{2},&\text{if $d\left(x,y\right)=2$};\\ 0,&\text{if $d\left(x,y\right)\geq 3$},\end{cases}$
from the distance-regularity. ∎
Next, we show two interesting properties of eigenvalues.
Lemma 6.
Let $G$ be a distance-regular graph with classical parameters $(3,b,\alpha,\beta)$ and Laplacian eigenvalues $0<\lambda_{1}<\lambda_{2}<\lambda_{3}$. Then we have
$\lambda_{1}+\lambda_{2}-\lambda_{3}\geq 0.$
Proof.
Recall the eigenvalues $\gamma_{1},\gamma_{2},\gamma_{3}$ from (16). We check directly that
$\gamma_{3}-\gamma_{1}$ $=(\beta+b-\alpha)(b+1),$
$\gamma_{3}-\gamma_{2}$ $=\beta+b^{2}-\alpha(b+1),$ (22)
$\gamma_{2}-\gamma_{1}$ $=b(\beta+1).$
Recall that $b$ is an integer not equal to 00 or $-1$. Applying $\beta\geq 1+(1+b)\alpha\geq 1$ from Lemma 4, we deduce that
$\begin{cases}0<\gamma_{1}<\gamma_{2}<\gamma_{3},&\text{if $b\geq 1$;}\\ 0<\gamma_{2}<\gamma_{3}<\gamma_{1},&\text{if $b\leq-2$.}\end{cases}$ (23)
If $b\geq 1$, we know $\beta\geq 1$ and $\alpha\geq 0$ from Lemma 4. Hence, we derive that
$\lambda_{1}+\lambda_{2}-\lambda_{3}=\gamma_{1}+\gamma_{2}-\gamma_{3}=2\alpha(1+b)+(b^{2}-1)(\beta-1)\geq 0.$
If $b\leq-2$, we again have $\beta\geq 1$ and hence
$\lambda_{1}+\lambda_{2}-\lambda_{3}=\gamma_{3}+\gamma_{2}-\gamma_{1}=(b+1)^{2}(\beta+1)\geq 0.$
This completes the proof. ∎
Lemma 7.
Let $G$ be a distance-regular graph with classical parameters $(3,b,\alpha,\beta)$ and Laplacian eigenvalues $0<\lambda_{1}<\lambda_{2}<\lambda_{3}$. Then for any two adjacent vertices $u$ and $v$ we have
$L^{2}(u,v)-L(u,v)(\lambda_{2}+\lambda_{3})\geq 0.$
Proof.
If $b\geq 1$, then we derive by (13), (16), (23) and Lemma 4 that
$L^{2}(u,v)-L(u,v)(\lambda_{2}+\lambda_{3})=-2d+a_{1}+(\gamma_{2}+\gamma_{3})$
$=$ $-2(1+b+b^{2})\beta+(-1+b\alpha+b^{2}\alpha+\beta)+(1+b)(1+\alpha+b\beta)+(1+b+b^{2})(1+\beta)$
$=$ $(1+b)^{2}(1+\alpha)>0.$
If $b\leq-2$, we deduce similarly that
$L^{2}(u,v)-L(u,v)(\lambda_{2}+\lambda_{3})=-2d+a_{1}+(\gamma_{3}+\gamma_{1})$
$=$ $-2(1+b+b^{2})\beta+(-1+b\alpha+b^{2}\alpha+\beta)+(1+b+b^{2})(1+\beta)+(1+\alpha+b\alpha+b^{2}\beta)$
$=$ $-b\beta+(1+b)^{2}\alpha+1+b+b^{2}$
$>$ $-b(1+(1+b)\alpha)+(1+b)^{2}\alpha+1+b+b^{2}$
$=$ $(1+b)\alpha+1+b^{2}$
$>$ $-(1+b)+1+b^{2}$
$=$ $b^{2}-b>0.$
This completes the proof. ∎
In the following two subsections, we present the proof of Theorem 1, in case of $b\geq 1$ and $b\leq-2$, respectively.
4.1 Proof of Theorem 1 in case $b\geq 1$
Let $G$ be a distance-regular graph with classical parameters $(3,b,\alpha,\beta)$ and Laplacian eigenvalues $0<\lambda_{1}<\lambda_{2}<\lambda_{3}$. We assume $b\geq 1$ in this subsection. By (23), we have
$\lambda_{1}=1+\alpha+b\alpha+b^{2}\beta,\lambda_{2}=(1+b)(1+\alpha+b\beta),\,\text{and}\,\,\lambda_{3}=(1+b+b^{2})(1+\beta).$ (24)
Lemma 8.
Let $G$ be a distance-regular graph with classical parameters $(3,b,\alpha,\beta)$, where $b\geq 1$. Then we have for any two distinct vertices $u$ and $v$ that
$\Delta_{1}(u,v)\geq 0,\Delta_{2}(u,v)\geq 0,and\ \Delta_{3}(u,v)\geq 0.$
Proof.
Inserting the expressions of $L^{2}(u,v)$ in Proposition 2 into (7)-(9) yields for any $i\in\{1,2,3\}$ that
$\Delta_{i}(u,v)=\begin{cases}C_{i}\left\{(\lambda_{j}-d-1)(\lambda_{k}-d-1)+d-a_{1}-1\right\},&\text{if $d(u,v)=1$;}\\ C_{i}\left\{(\lambda_{j}-d)(\lambda_{k}-d)+d-c_{2}\right\},&\text{if $d(u,v)=2$;}\\ C_{i}\left\{(\lambda_{j}-d)(\lambda_{k}-d)+d\right\},&\text{if $d(u,v)=3$,}\end{cases}$ (25)
where $\{j,k\}=\{1,2,3\}\setminus\{i\}$.
Recalling that $C_{1}>0$, $C_{2}<0$, $C_{3}>0$, and $c_{2}>0$, it remains to show
  • (i)
    $(\lambda_{2}-d-1)(\lambda_{3}-d-1)+d-a_{1}-1\geq 0$, and $(\lambda_{2}-d)(\lambda_{3}-d)+d-c_{2}\geq 0$.
  • (ii)
    $(\lambda_{1}-d-1)(\lambda_{3}-d-1)+d-a_{1}-1\leq 0$, and $(\lambda_{1}-d)(\lambda_{3}-d)+d\leq 0$.
  • (iii)
    $(\lambda_{1}-d-1)(\lambda_{2}-d-1)+d-a_{1}-1\geq 0$, and $(\lambda_{1}-d)(\lambda_{2}-d)+d-c_{2}\geq 0$.
Notice that $d-\lambda_{i}=d-\gamma_{i}=\theta_{i}$ whose expressions are provided in (15). Next, we check the above inequalities case by case.
Case $(i)$: Inserting the expression $d-a_{1}-1=b_{1}=(b+b^{2})(\beta-\alpha)$ from (13), we deduce
$(\lambda_{2}-d-1)(\lambda_{3}-d-1)+d-a_{1}-1=(1+\theta_{2})(1+\theta_{3})+b_{1}$
$=$ $\left\{1+\beta-(1+b)(1+\alpha)\right\}(-b-b^{2})+(b+b^{2})(\beta-\alpha)$
$=$ $b^{2}(1+b)(1+\alpha)>0,$ (26)
where the last inequality follows from the fact that $\alpha\geq 0$ due to Lemma 4. Furthermore, we derive by noticing $d=(1+b+b^{2})\beta$ and $c_{2}=(1+b)(1+\alpha)$ that
$(\lambda_{2}-d)(\lambda_{3}-d)+d-c_{2}=\theta_{2}\theta_{3}+d-c_{2}$
$=$ $-\left\{\beta-(1+b)(1+\alpha)\right\}(1+b+b^{2})+(1+b+b^{2})\beta-c_{2}$
$=$ $c_{2}(b+b^{2})>0.$ (27)
Case $(ii)$: Similarly as in Case $(i)$, we derive
$(\lambda_{1}-d-1)(\lambda_{3}-d-1)+d-a_{1}-1=(1+\theta_{1})(1+\theta_{3})+b_{1}$
$=$ $(1+b)(\beta-\alpha)(-b-b^{2})+(b+b^{2})(\beta-\alpha)$
$=$ $-(b+b^{2})b(\beta-\alpha)<0,$ (28)
and
$(\lambda_{1}-d)(\lambda_{3}-d)+d=\theta_{1}\theta_{3}+d$
$=$ $-\left\{(1+b)(\beta-\alpha)-1\right\}(1+b+b^{2})+(1+b+b^{2})\beta$
$=$ $(1+b+b^{2})(1+(1+b)\alpha-b\beta)\leq 0,$ (29)
where we have used $\beta\geq 1+(1+b)\alpha$.
Case $(iii)$: Similarly as in Case $(i)$, we deduce
$(\lambda_{1}-d-1)(\lambda_{2}-d-1)+d-a_{1}-1=(1+\theta_{1})(1+\theta_{2})+b_{1}$
$=$ $(1+b)(\beta-\alpha)\left\{1+\beta-(1+b)(1+\alpha)\right\}+(b+b^{2})(\beta-\alpha)$
$=$ $(1+b)(\beta-\alpha)(\beta-b\alpha-\alpha)\geq 0,$ (30)
and
$(\lambda_{1}-d)(\lambda_{2}-d)+d-c_{2}=\theta_{1}\theta_{2}+d-c_{2}$
$=$ $\left\{(1+b)(\beta-\alpha)-1\right\}\left\{\beta-(1+b)(1+\alpha)\right\}+(1+b+b^{2})\beta-c_{2}$
$=$ $(1+b)(\beta-\alpha-1)(\beta-b\alpha-\alpha)\geq 0,$ (31)
where we have used $\beta\geq 1+(1+b)\alpha$ and $\alpha\geq 0$. This completes the proof. ∎
Lemma 9.
Let $G$ be a distance-regular graph with classical parameters $(3,b,\alpha,\beta)$, where $b\geq 1$. For any two distinct vertices $u$ and $v$, we have the following estimates.
  • (i)
    If $d(u,v)=1$, then $\Delta_{12}(u,v)\geq 0$.
  • (ii)
    If $d(u,v)=2$ and $\beta\geq 1+(2+b)\alpha$, then $\Delta_{12}(u,v)\geq 0$.
  • (iii)
    If $d(u,v)=2$ and $\beta<1+(2+b)\alpha$, then $\Delta_{13}(u,v)\leq 0$.
  • (iv)
    If $d(u,v)=3$, then $\Delta_{23}(u,v)\geq 0$.
Proof.
For $\{i,j,k\}=\{1,2,3\}$, we obtain by rearranging (10) that
$\frac{n\Delta_{ij}(u,v)}{C_{i}C_{j}(\lambda_{j}-\lambda_{i})}=(d+(n-1)L(u,v))\lambda_{k}^{2}-$ $\left((n-1)L^{2}(u,v)+d^{2}+d\right)\lambda_{k}$
$+$ $nd\left(L^{2}(u,v)-(d+1)L(u,v)\right).$
Applying Proposition 2, we further deduce that
$\frac{n\Delta_{ij}(u,v)}{C_{i}C_{j}(\lambda_{j}-\lambda_{i})}=\begin{cases}d\lambda_{k}(\lambda_{k}-d-1),&\text{if $d(u,v)=3$;}\\ d\lambda_{k}(\lambda_{k}-d-1)-\left((n-1)\lambda_{k}-nd\right)c_{2},&\text{if $d(u,v)=2$;}\\ -(n-1-d)\lambda_{k}^{2}+\left((n-1)(2d-a_{1})-d(d+1)\right)\lambda_{k}&\\ \hskip 156.49014pt-nd(d-a_{1}-1),&\text{if $d(u,v)=1$.}\end{cases}$
Inserting $d-a_{1}-1=b_{1}$, we have in the case $d(u,v)=1$ that
$\frac{n\Delta_{ij}(u,v)}{C_{i}C_{j}(\lambda_{j}-\lambda_{i})}=$ $-(n-1-d)\lambda_{k}^{2}+\left((n-1-d)(d+1)+(n-1)b_{1}\right)\lambda_{k}-ndb_{1}$
$=$ $-(n-1-d)\lambda_{k}(\lambda_{k}-d-1)+\left((n-1)\lambda_{k}-nd\right)b_{1}.$ (32)
Case $(i)$: $d(u,v)=1$. Recall that $\lambda_{3}=\gamma_{3}=(1+b+b^{2})(1+\beta)=d+1+b+b^{2}$. We deduce
$\frac{n\Delta_{12}(u,v)}{C_{1}C_{2}(\lambda_{2}-\lambda_{1})}=$ $-(n-1-d)(d+1+b+b^{2})(b+b^{2})+\left((n-1)(1+b+b^{2})-d\right)b_{1}$
$=$ $(1+b+b^{2})\left\{-(n-1-d)(1+\beta)b(1+b)+(n-1-\beta)b_{1}\right\}$
Using $b_{1}=(b+b^{2})(\beta-\alpha)$, we arrive at
$\frac{n\Delta_{12}(u,v)}{C_{1}C_{2}(\lambda_{2}-\lambda_{1})}=$ $(1+b+b^{2})b(1+b)\left\{-(n-1-d)(1+\beta)+(n-1-\beta)(\beta-\alpha)\right\}$
$=$ $(1+b+b^{2})b(1+b)\left\{-(n-1)(1+\alpha)+d(1+\beta)-\beta(\beta-\alpha)\right\}.$ (33)
By (2.4), we have
$(n-1)(1+\alpha)=d(1+\alpha)+db(\beta-\alpha)+db(\beta-\alpha)\frac{b_{2}}{c_{3}}.$ (34)
Therefore, we deduce
$-(n-1)(1+\alpha)+d(1+\beta)-\beta(\beta-\alpha)$
$<$ $-d(1+\alpha)-db(\beta-\alpha)+d(1+\beta)-\beta(\beta-\alpha)$
$=$ $(\beta-\alpha)(-db+b-\beta)<0.$
Inserting into to (33), we obtain
$\frac{n\Delta_{12}(u,v)}{C_{1}C_{2}(\lambda_{2}-\lambda_{1})}<0,$
which implies that $\Delta_{12}(u,v)>0$.
Case $(ii)$: $d(u,v)=2$ and $\beta\geq 1+(2+b)\alpha$. We have in this case that
$\frac{n\Delta_{12}(u,v)}{C_{1}C_{2}(\lambda_{2}-\lambda_{1})}=$ $d\lambda_{3}(\lambda_{3}-d-1)-\left((n-1)\lambda_{3}-nd\right)c_{2}=:B_{1}+B_{2}.$ (35)
Recall that $\lambda_{3}=(1+b+b^{2})(1+\beta)$, $d=(1+b+b^{2})\beta$ and $c_{2}=(1+b)(1+\alpha)$. We deduce that
$B_{1}:=d\lambda_{3}(\lambda_{3}-d-1)=(1+b+b^{2})db(1+b)(1+\beta),$
and
$B_{2}:=-\left((n-1)\lambda_{3}-nd\right)c_{2}=-(1+b+b^{2})(1+b)(1+\alpha)(n-1-\beta).$
Applying (2.4), we further calculate that
$\frac{B_{1}+B_{2}}{(1+b)(1+b+b^{2})}=-(n-1-\beta)(1+\alpha)+db(1+\beta)$
$=$ $-(d-\beta)(1+\alpha)-db(\beta-\alpha)-db(\beta-\alpha)\frac{b_{2}}{c_{3}}+db(1+\beta)$
$=$ $(db-d+\beta)(1+\alpha)-db(\beta-\alpha)\frac{b_{2}}{c_{3}}$
$=$ $b\beta\left\{b^{2}-(1+b+b^{2})(\beta-\alpha)\frac{b_{2}}{c_{3}}.\right\}$
Inserting the expressions $b_{2}=b^{2}(\beta-\alpha-b\alpha)$ and $c_{3}=(1+b+b^{2})(1+\alpha+b\alpha)$, we obtain
$\frac{(B_{1}+B_{2})(1+\alpha+b\alpha)}{b(1+b)(1+b+b^{2})\beta}=$ $b^{2}(1+\alpha+b\alpha)-(\beta-\alpha)b^{2}(\beta-\alpha-b\alpha)$
$=$ $-b^{2}(1+\beta)(\beta-1-(2+b)\alpha).$
By assumption, we have $\beta-1-(2+b)\alpha\geq 0$, and hence $B_{1}+B_{2}\leq 0$. Hence, we derive $\Delta_{12}(u,v)\geq 0$ by (35).
Case $(iii)$: $d(u,v)=2$ and $\beta<1+(2+b)\alpha$. We have in this case that
$\frac{n\Delta_{13}(u,v)}{C_{1}C_{3}(\lambda_{3}-\lambda_{1})}=$ $d\lambda_{2}(\lambda_{2}-d-1)-\left((n-1)\lambda_{2}-nd\right)c_{2}$
$=$ $\left(d\lambda_{2}-(n-1)c_{2}\right)(\lambda_{2}-d-1)-(n-1-d)c_{2}$
$=$ $(1+b)\left\{d(1+\alpha+b\beta)-(n-1)(1+\alpha)\right\}(b-\beta+b\alpha+\alpha)$
$\hskip 142.26378pt-(n-1-d)(1+b)(1+\alpha),$
where we have used $\lambda_{2}=(1+b)(1+\alpha+b\beta)$, $d=(1+b+b^{2})\beta$ and $c_{2}=(1+b)(1+\alpha)$. Applying (2.4), we derive
$\frac{n\Delta_{13}(u,v)}{C_{1}C_{3}(\lambda_{3}-\lambda_{1})(1+b)}$
$=$ $\left(db\alpha-db(\beta-\alpha)\frac{b_{2}}{c_{3}}\right)(b-\beta+b\alpha+\alpha)-\left(db(\beta-\alpha)+db(\beta-\alpha)\frac{b_{2}}{c_{3}}\right)$
$=$ $db\left\{\alpha(b-\beta+b\alpha+\alpha)-(\beta-\alpha)\right\}-db(\beta-\alpha)\frac{b_{2}}{c_{3}}(b-\beta+b\alpha+\alpha+1)$
$=$ $-db(1+\alpha)(\beta-\alpha-b\alpha)-\beta b^{3}(\beta-\alpha)\frac{\beta-\alpha-b\alpha}{1+\alpha+b\alpha}(b-\beta+b\alpha+\alpha+1),$
where we have inserted $b_{2}=b^{2}(\beta-\alpha-b\alpha)$, $c_{3}=(1+b+b^{2})(1+\alpha+b\alpha)$ and $d=(1+b+b^{2})\beta$. Rearranging the identity yields
$\frac{n(1+\alpha+b\alpha)\Delta_{13}(u,v)}{C_{1}C_{3}(\lambda_{3}-\lambda_{1})\beta b(1+b)(\beta-\alpha-b\alpha)}$
$=$ $-(1+b+b^{2})(1+\alpha)(1+\alpha+b\alpha)+b^{2}(\beta-\alpha)(\beta-b-b\alpha-\alpha-1)=:g(\beta).$ (36)
For the function $g(\beta)$, we observe that
$\max_{1+(1+b)\alpha\leq\beta<1+(2+b)\alpha}g(\beta)\leq\max\{g(1+(1+b)\alpha),g(1+(2+b)\alpha)\}.$
Since $-(1+b+b^{2})(1+\alpha)(1+\alpha+b\alpha)<0$, and the value $1+(1+b)\alpha$ lies in between $\alpha$ and $b+b\alpha+\alpha+1$, we have $g(1+(1+b)\alpha)<0$. Moreover, we check
$g(1+(2+b)\alpha)=(1+\alpha+b\alpha)(-b^{3}-(1+b+b^{2})-(1+b)(1+\alpha))<0.$
Here, we used $\alpha\geq 0$ from Lemma 4. Therefore, we have $\max_{1+(1+b)\alpha\leq\beta<1+(2+b)\alpha}g(\beta)<0$, and hence $\Delta_{13}(u,v)>0$ by (4.1).
Case $(iv)$: $d(u,v)=3$. Recall that $d-\lambda_{1}=\theta_{1}=(1+b)(\beta-\alpha)-1$. We derive that
$\frac{n\Delta_{23}(u,v)}{C_{2}C_{3}(\lambda_{3}-\lambda_{2})}=d\lambda_{1}(\lambda_{1}-d-1)=-d\lambda_{1}(1+b)(\beta-\alpha)<0.$
Indeed, we have for any graph with $n$ vertices and maximal degree $d_{\max}$, we have $\lambda_{1}\leq\frac{n}{n-1}d_{\max}<d_{\max}$ by a trace argument. Hence we have $\lambda_{1}-d<0$ for any $d$-regular graph. Therefore, we obtain $\Delta_{23}(u,v)>0$. This completes the proof. ∎
Proof of Theorem 1 in case $b\geq 1$.
In case $b\geq 1$, Theorem 1 follows from Theorem 2, by verifing the conditions via Lemma 6, Lemma 7, Lemma 8 and Lemma 9. ∎
4.2 Proof of Theorem 1 in case $b\leq-2$
In this section , we will prove Theorem 1 in the case of $b\leq-2$.
Lemma 10.
Let $G$ be a distance-regular graph with classical parameters $(3,b,\alpha,\beta)$, where $b\leq-2$. Then we have for any two distinct vertices $u$ and $v$ that
$\Delta_{1}(u,v)\geq 0,\Delta_{2}(u,v)\geq 0,\text{and}\ \Delta_{3}(u,v)\geq 0.$
Proof.
The proof is similar to that of Lemma 8. Recalling that $C_{1}>0$, $C_{2}<0$, $C_{3}>0$, and $c_{2}\geq 0$ and the identity (25), it remains to show
  • (i)
    $(\lambda_{2}-d-1)(\lambda_{3}-d-1)+d-a_{1}-1\geq 0$, and $(\lambda_{2}-d)(\lambda_{3}-d)+d-c_{2}\geq 0$.
  • (ii)
    $(\lambda_{1}-d-1)(\lambda_{3}-d-1)+d-a_{1}-1\leq 0$, and $(\lambda_{1}-d)(\lambda_{3}-d)+d\leq 0$.
  • (iii)
    $(\lambda_{1}-d-1)(\lambda_{2}-d-1)+d-a_{1}-1\geq 0$, and $(\lambda_{1}-d)(\lambda_{2}-d)+d-c_{2}\geq 0$.
By (23), we have in case $b\leq-2$ that $\lambda_{1}=\gamma_{2},\lambda_{2}=\gamma_{3}$ and $\lambda_{3}=\gamma_{1}$. This tells in particular that
$d-\lambda_{1}=\theta_{2},\,\,d-\lambda_{2}=\theta_{3},\,\,\text{and}\,\,d-\lambda_{3}=\theta_{1}.$
Using the calculations in the proof of Lemma 8, we check $(i)-(iii)$ as follows.
Case $(i)$: By (28), we have
$(\lambda_{2}-d-1)(\lambda_{3}-d-1)+(d-a_{1}-1)=(1+\theta_{3})(1+\theta_{1})+b_{1}$
$=$ $-b(b+b^{2})(\beta-\alpha)>0.$
In the last inequality, we have used the fact that $\alpha<-1$ due to Lemma 4. Furthermore, we derive by (29) that
$(\lambda_{2}-d)(\lambda_{3}-d)+d-c_{2}=\theta_{3}\theta_{1}+d-c_{2}$
$=$ $(1+b+b^{2})(1+(1+b)\alpha-b\beta)-(1+b)(1+\alpha)$
$=$ $b\left\{b+(1+b)^{2}\alpha-(1+b+b^{2})\beta\right\}>0.$
Case $(ii)$: We derive by (30) that
$(\lambda_{1}-d-1)(\lambda_{3}-d-1)+d-a_{1}-1=(1+\theta_{2})(1+\theta_{1})+b_{1}$
$=$ $(1+b)(\beta-\alpha)(\beta-b\alpha-\alpha)<0.$
Furthermore, we deduce from (31) that
$(\lambda_{1}-d)(\lambda_{3}-d)+d=\theta_{2}\theta_{1}+d$
$=$ $(1+b)(\beta-\alpha-1)(\beta-b\alpha-\alpha)+c_{2}$
$=$ $(1+b)\left\{\beta(\beta-b\alpha-\alpha)-(1+\alpha)(\beta-b\alpha-\alpha-1)\right\}<0.$
Case $(iii)$: We deduce from (26) and (27) that
$(\lambda_{1}-d-1)(\lambda_{2}-d-1)+d-a_{1}-1=(1+\theta_{2})(1+\theta_{3})+b_{1}$
$=$ $b^{2}(1+b)(1+\alpha)>0,$
and
$(\lambda_{1}-d)(\lambda_{2}-d)+d-c_{2}=\theta_{2}\theta_{3}+d-c_{2}=c_{2}(b+b^{2})>0.$
This completes the proof. ∎
Lemma 11.
Let $G$ be a distance-regular graph with classical parameters $(3,b,\alpha,\beta)$, where $b\leq-2$. For any two distinct vertices $u$ and $v$, we have the following estimates.
  • (i)
    If $d(u,v)=1$, then $\Delta_{12}(u,v)\geq 0$,
  • (ii)
    If $d(u,v)=2$, then $\Delta_{12}(u,v)\geq 0$,
  • (iii)
    If $d(u,v)=3$, then $\Delta_{23}(u,v)\geq 0$.
Proof.
As in the proof of Lemma 9, we have for $\{i,j,k\}=\{1,2,3\}$ that
$\frac{n\Delta_{ij}(u,v)}{C_{i}C_{j}(\lambda_{j}-\lambda_{i})}=\begin{cases}d\lambda_{k}(\lambda_{k}-d-1),&\text{if $d(u,v)=3$;}\\ d\lambda_{k}(\lambda_{k}-d-1)-\left((n-1)\lambda_{k}-nd\right)c_{2},&\text{if $d(u,v)=2$;}\\ -(n-1-d)\lambda_{k}(\lambda_{k}-d-1)+\left((n-1)\lambda_{k}-nd\right)b_{1},&\text{if $d(u,v)=1$.}\end{cases}$
Next, we check the three cases in reverse order.
Case $(iii)$: $d(u,v)=3$. We have
$\frac{n\Delta_{23}(u,v)}{C_{2}C_{3}(\lambda_{3}-\lambda_{2})}=d\lambda_{1}(\lambda_{1}-d-1)<0.$
The inequality follows from the general fact that $\lambda_{1}<d$ for any $d$-regular graph, which can be shown by a trace argument. The inequality can also be checked by inserting the expressions of $\lambda_{1}=\gamma_{2}$ and $d$. Therefore, we have $\Delta_{23}(u,v)>0$.
Case $(ii)$: $d(u,v)=2$. In this case we have
$\frac{n\Delta_{12}(u,v)}{C_{1}C_{2}(\lambda_{2}-\lambda_{1})}=$ $d\lambda_{3}(\lambda_{3}-d-1)-\left((n-1)\lambda_{3}-nd\right)c_{2}$
$=\left(d\lambda_{3}-(n-1)c_{2}\right)(\lambda_{3}-d-1)-(n-1-d)c_{2}.$
Recall that $c_{2}=(1+b)(1+\alpha)$, $\lambda_{3}=\gamma_{1}=1+\alpha+b\alpha+b^{2}\beta$, and $d=(1+b+b^{2})\beta$. We derive by (2.4) that
$\frac{n\Delta_{12}(u,v)}{C_{1}C_{2}(\lambda_{2}-\lambda_{1})}$
$=$ $\left(d\lambda_{3}-(n-1)(1+\alpha)(1+b)\right)(\alpha-\beta)(1+b)-(n-1-d)(1+\alpha)(1+b)$
$=$ $\left(d\lambda_{3}-d(1+\alpha)(1+b)-db(\beta-\alpha)\left(1+\frac{b_{2}}{c_{3}}\right)(1+b)\right)(\alpha-\beta)(1+b)$
$\hskip 199.16928pt-db(\beta-\alpha)\left(1+\frac{b_{2}}{c_{3}}\right)(1+b)$
$=$ $d(b^{2}\beta-b)(\alpha-\beta)(1+b)-db(\beta-\alpha)\left(1+\frac{b_{2}}{c_{3}}\right)(1+b)((\alpha-\beta)(1+b)+1).$
Recall that
$d\left(1+\frac{b_{2}}{c_{3}}\right)=\beta\frac{(1+\alpha)(1+b+b^{2})+b(\alpha+b\beta)}{1+\alpha+b\alpha}.$
Then, we continue to calculate
$\frac{n\Delta_{12}(u,v)}{C_{1}C_{2}(\lambda_{2}-\lambda_{1})}=\frac{b\beta(\alpha-\beta)(1+b)}{1+\alpha+b\alpha}\big{\{}(1+b+b^{2})(b\beta-1)(1+\alpha+b\alpha)$
$\hskip 99.58464pt+\left((1+\alpha)(1+b+b^{2})+b(\alpha+b\beta)\right)((\alpha-\beta)(1+b)+1)\big{\}}.$ (37)
By the rearrangement $(\alpha-\beta)(1+b)-1=\alpha+b\alpha-\beta-(b\beta-1)$, we deduce
$(1+b+b^{2})(b\beta-1)(1+\alpha+b\alpha)+\left((1+\alpha)(1+b+b^{2})+b(\alpha+b\beta)\right)((\alpha-\beta)(1+b)+1)$
$=$ $b^{2}\left(\alpha+b\alpha-\beta\right)(b\beta-1)+\left\{(1+\alpha)(1+b+b^{2})+b(\alpha+b\beta)\right\}(\alpha+b\alpha-\beta)$
$=$ $(\alpha+b\alpha-\beta)(1+b)(1+\alpha+b\alpha+b^{2}\beta).$
Inserting into (37) yields
$\frac{n\Delta_{12}(u,v)}{C_{1}C_{2}(\lambda_{2}-\lambda_{1})}=$ $d\lambda_{3}(\lambda_{3}-d-1)-\left((n-1)\lambda_{3}-nd\right)c_{2}$
$=$ $\frac{b\beta(\alpha-\beta)(1+b)}{1+\alpha+b\alpha}(\alpha+b\alpha-\beta)(1+b)(1+\alpha+b\alpha+b^{2}\beta).$ (38)
This implies that $\Delta_{12}(u,v)>0$.
Case $(i)$: $d(u,v)=1$. We have in this case that
$\frac{-c_{2}}{b_{1}}\cdot\frac{n\Delta_{12}(u,v)}{C_{1}C_{2}(\lambda_{2}-\lambda_{1})}=$ $(n-1-d)\lambda_{3}(\lambda_{3}-d-1)\frac{1+\alpha}{b(\beta-\alpha)}-\left((n-1)\lambda_{3}-nd\right)c_{2}$
$=$ $d\left(1+\frac{b_{2}}{c_{3}}\right)\lambda_{3}(\lambda_{3}-d-1)-\left((n-1)\lambda_{3}-nd\right)c_{2}.$
In the above, we have used $b_{1}=(b+b^{2})(\beta-\alpha)$, $c_{2}=(1+b)(1+\alpha)$ and (2.4). Using (38), we obtain
$\frac{-c_{2}}{b_{1}}\cdot\frac{n\Delta_{12}(u,v)}{C_{1}C_{2}(\lambda_{2}-\lambda_{1})}$
$=$ $\frac{b\beta(\alpha-\beta)(1+b)}{1+\alpha+b\alpha}(\alpha+b\alpha-\beta)(1+b)(1+\alpha+b\alpha+b^{2}\beta)+d\lambda_{3}(\lambda_{3}-d-1)\frac{b_{2}}{c_{3}}.$
Observing that
$d\lambda_{3}(\lambda_{3}-d-1)\frac{b_{2}}{c_{3}}=\frac{b\beta(\alpha-\beta)(1+b)}{1+\alpha+b\alpha}b(\beta-\alpha-b\alpha)(1+\alpha+b\alpha+b^{2}\beta)$
Hence, we derive
$\frac{-c_{2}}{b_{1}}\cdot\frac{n\Delta_{12}(u,v)}{C_{1}C_{2}(\lambda_{2}-\lambda_{1})}=$ $\frac{b\beta(\alpha-\beta)(1+b)}{1+\alpha+b\alpha}(\alpha+b\alpha-\beta)(1+\alpha+b\alpha+b^{2}\beta)>0.$
Therefore, we have $\Delta_{12}(u,v)>0$. This completes the proof. ∎
Proof of Theorem 1 in case $b\leq-2$.
Combining Lemma 6, Lemma 7, Lemma 10, Lemma 11 and Theorem 2, we prove Theorem 1 in case $b\leq-2$. ∎
Remark 5.
Known distance regular graphs with classical parameters of diameter at least $3$ are listed in [3, Tables 6.1 and 6.2]. For classification results of distance regular graphs with classical parameters of diameter at least $4$ and $b\leq-2$ or $b=1$, we refer to [3, 6, 13, 14] and the references therein. It is natural to expect the monotonic normalized heat diffusion property holds for any distance regular graphs with classical parameters.
5 Further Discussion
The distance-regular graphs with diameter $3$ are classified as antipodal, bipartite, and primitive [2]. A distance-regular graph with an intersection array $\{d,b_{1},b_{2};1,c_{2},c_{3}\}$ is antipodal if $b_{2}=1$ and $c_{3}=d$. In fact, the general form of the intersection array for an antipodal distance-regular graph of diameter $3$ is $\{b_{0},b_{1},b_{2};c_{1},c_{2},c_{3}\}=\{d,m\gamma,1;1,\gamma,d\}$, where $m$ is a positive integer [2, Theorem 2]. A distance-regular graph of diameter $3$ is primitive if it is neither antipodal nor bipartite.
For the bipartite case, the MNHD property is true by the result of Kubo and Namba [8]. For the antipodal case, we show that the MNHD property holds still.
Theorem 3.
The antipodal distance-regular graphs with diameter $3$ have the MNHD property.
Proof.
Let $G$ be an antipodal distance-regular graph with intersection array $\{d,m\gamma,1;1,\gamma,d\}$, where $m$ is a positive integer. Let us denote $M:=\sqrt{4d+(d-\gamma-m\gamma-1)^{2}}$. The non-trivial Laplacian eigenvalues are listed below,
$\lambda_{1}$ $=d+\frac{1}{2}(1-d+\gamma+m\gamma-M),$
$\lambda_{2}$ $=1+d,$
$\lambda_{3}$ $=d+\frac{1}{2}(1-d+\gamma+m\gamma+M),$
with $\lambda_{1}<\lambda_{2}<\lambda_{3}$. The adjacency eigenvalues of an antipodal distance-regular graph of diameter $3$ have been calculated in [2, Page 75]. One can also derive them directly from the intersection matrix as explained in [6, Page 13].
We further check that $n=1+d+m+dm$.
Next, we check the conditions in Theorem 2. First, we observe that $\lambda_{1}+\lambda_{2}-\lambda_{3}\geq 0$. Indeed, we have $\lambda_{1}+\lambda_{2}-\lambda_{3}=1+d-M$. As $(1+d)^{2}-M^{2}=\gamma(1+m)(2d-m-m\gamma-2)\geq 0$, we derive $1+d-M\geq 0$.
Moreover, we have $L^{2}(u,v)-L(u,v)(\lambda_{2}+\lambda_{3})\geq 0$ for any two adjacent vertices $u$ and $v$. Recall that $L^{(}u,v)=-2d+a_{1}=-2d+(d-1-m\gamma)$ and $L(u,v)=-1$. We get $L^{2}(u,v)-L(u,v)(\lambda_{2}+\lambda_{3})=\frac{1}{2}(1+d+\gamma-m\gamma+M)\geq 0$, as desired.
Case $(i)$: $d(u,v)=1$. It follows from the identity (25) that
$\Delta_{1}(u,v)/C_{1}=(\lambda_{2}-d-1)(\lambda_{3}-d-1)+d-a_{1}-1.$
Noting that $\lambda_{2}=1+d$ and $d-a_{1}-1=b_{1}=m\gamma$, we deduce $\Delta_{1}/C_{1}=m\gamma\geq 0$. Since $C_{1}>0$, we obtain $\Delta_{1}(u,v)\geq 0$.
Similarly, we have
$\Delta_{3}(u,v)/C_{3}=$ $(\lambda_{1}-d-1)(\lambda_{2}-d-1)+d-a_{1}-1=m\gamma>0.$
This implies $\Delta_{3}(u,v)\geq 0$. For $\Delta_{2}(u,v)$, we calclulate
$\Delta_{2}(u,v)/C_{2}=$ $(\lambda_{1}-d-1)(\lambda_{3}-d-1)+d-a_{1}-1$
$=$ $\frac{1}{2}(-1-d+\gamma+m\gamma-M)\cdot\frac{1}{2}(-1-d+\gamma+m\gamma+M)+m\gamma$
$=$ $-\gamma<0.$
Since $C_{2}<0$, we obtain $\Delta_{2}(u,v)>0$.
By (32), we have
$\frac{n\Delta_{12}(u,v)}{C_{1}C_{2}(\lambda_{2}-\lambda_{1})}=$ $-(n-1-d)\lambda_{3}(\lambda_{3}-d-1)+\left((n-1)\lambda_{3}-nd\right)b_{1}$
$=$ $-(n-1-d)\lambda_{3}(\lambda_{3}-d-1)+n(\lambda_{3}-d)b_{1}-\lambda_{3}b_{1}$
$=$ $-\lambda_{3}b_{1}<0,$
where we used
$-(n-1-d)\lambda_{3}(\lambda_{3}-d-1)+n(\lambda_{3}-d)b_{1}$
$=$ $-m(d+1)\frac{1}{4}(1+d+\gamma+m\gamma+M)(-1-d+\gamma+m\gamma+M)$
$\hskip 85.35826pt+\frac{m\gamma}{2}(d+1)(m+1)(1-d+\gamma+m\gamma+M)$
$=$ $-\frac{m(d+1)}{4}\left\{(M+\gamma(1+m))^{2}-(d+1)^{2}-2\gamma(m+1)(1-d+\gamma(m+1)+M)\right\}$
$=$ $0$
Now we apply Theorem 2 to conclude that the function $r_{t}(u,v):=\frac{H_{t}(u,v)}{H_{t}(u,u)}$ is monotonically non-decreasing in $t$ for any two adjacent vertices $u$ and $v$.
Case $(ii)$: $d(u,v)=2$. Similar as in Case (i), we have
$\Delta_{1}(u,v)/C_{1}=$ $(\lambda_{2}-d)(\lambda_{3}-d)+d-c_{2}$
$=$ $\frac{1}{2}(1-d+\gamma+m\gamma+M)+d-\gamma$
$=$ $\frac{1}{2}\left(1+d-\gamma+m\gamma+M\right)>0,$
$\Delta_{2}(u,v)/C_{2}=(\lambda_{1}-d)(\lambda_{3}-d)+d-c_{2}=-\gamma<0.$
and
$\Delta_{3}(u,v)/C_{3}=(\lambda_{1}-d)(\lambda_{2}-d)+d-c_{2}=\frac{1}{2}\left(1+d-\gamma+m\gamma-M\right)>0,$
where in the last inequlaity we used $(1+d-\gamma+m\gamma)^{2}-M^{2}=-4\gamma(1-dm+m\gamma)>0$. Therefore, we obtain $\Delta_{i}(u,v)\geq 0$ for each $i=1,2,3$.
Furthermore, we check
$\frac{n\Delta_{13}(u,v)}{C_{1}C_{3}(\lambda_{3}-\lambda_{1})}=d\lambda_{2}(\lambda_{2}-d-1)-\left((n-1)\lambda_{2}-nd\right)c_{2}=-(1+d)m\gamma<0,$
and, hence, $\Delta_{13}(u,v)<0$. We apply Theorem 2 to conclude that the function $r_{t}(u,v)$ is monotonically non-decreasing in $t$ for any two vertices $u$ and $v$ with $d(u,v)=2.$
Case $(iii)$: $d(u,v)=2$. Similar as in Case (i), we have
$\Delta_{1}(u,v)/C_{1}=(\lambda_{2}-d)(\lambda_{3}-d)+d=\frac{1}{2}\left(1+d+\gamma+m\gamma+M\right)>0,$
$\Delta_{2}(u,v)/C_{2}=(\lambda_{1}-d)(\lambda_{3}-d)+d=0,$
and
$\Delta_{3}(u,v)/C_{3}=(\lambda_{1}-d)(\lambda_{2}-d)+d=\frac{1}{2}\left(1+d+\gamma+m\gamma-M\right)>0,$
where in the last inequality we used $(1+d+\gamma+m\gamma)^{2}-M^{2}=4d(1+m)\gamma>0$. Hence, we obtain $\Delta_{i}(u,v)\geq 0$ for each $i=1,2,3$.
Furthermore, we check
$\frac{n\Delta_{13}(u,v)}{C_{1}C_{3}(\lambda_{3}-\lambda_{1})}=d\lambda_{2}(\lambda_{2}-d-1)=0.$
Applying Theorem 2, we derive that the function $r_{t}(u,v)$ is monotonically non-decreasing in $t$ for any two vertices $u$ and $v$ with $d(u,v)=3$. ∎
Our main Theorem 1 implies that the MNHD property holds for primitive distance-regular graphs with classical parameters of diameter $3$. It is natural to ask whether all distance-regular graphs with diameter $3$ have the MNHD property or not.
Acknowledgement
H.Z. is very grateful to Kaizhe Chen and Chenhui Lv for their helpful discussions. This work is supported by National Key R & D Program of China 2023YFA1010200 and National Natural Science Foundation of China No. 12031017 and No. 12431004.

### v2

@@
-Monotonic normalized heat diffusion for distance-regular graphs with classical parameters of diameter $3$
+Monotonic normalized heat diffusion for distance-regular graphs with classical parameters of diameter 33
 Abstract
@@
 However, there do exist graphs for which the function $r_{t}(u,v)$ exceeds 1 for certain vertices $u,v$ at certain time $t$.
-This is observed by Regev and Shinkar [12, Proposition A.1].
+This is observed by Regev and Shinkar [15, Proposition A.1].
 Regev and Shinkar mentioned that this observation was brought to their attention by Cheeger.
-They also provided an explicit example which is a $4$-regular graph with $10$ vertices [12, Figure 2].
-A sufficient condition to ensure that $r_{t}(u,v)\leq 1$ is a heat homogeneity, that is, the heat kernel has constant diagonal, which means that the value of $H_{t}(u,u)$ is independent of the choice of $u$ (see [12, Section 1] and [10, Section 1]).
+They also provided an explicit example which is a $4$-regular graph with $10$ vertices [15, Figure 2].
+A sufficient condition to ensure that $r_{t}(u,v)\leq 1$ is a heat homogeneity, that is, the heat kernel has constant diagonal, which means that the value of $H_{t}(u,u)$ is independent of the choice of $u$ (see [15, Section 1] and [10, Section 1]).
 In fact, the heat homogeneity holds if and only if the graph is walk-regular [10, Theorem A.1].
@@
 Typical walk-regular graphs includes vertex-transitive graphs, distance-regular graphs, and regular graphs with at most $4$ distinct eigenvalues.
-Peres asked in 2013 (see [12, Section 1]) whether the MNHD property holds for all vertex transitive graphs or not.
+Peres asked in 2013 (see [15, Section 1]) whether the MNHD property holds for all vertex transitive graphs or not.
 Regev and Shinkar [15] showed that there exists a Cayley graph for which the MNHD property fails.
@@
 Notice that the bipartite case follows from the result of Kubo and Namba [8].
+Relevance of the MNHD property in other areas
+It is very natural to study whether the MNHD property holds on a given graph.
+In general, it is highly concerned to understand how certain properties of solutions to differential/difference equations depend on the structure of the underlying graphs.
+Moreover, the MNHD property is very useful and relevant to other research areas.
+It is also very interesting to study the MNHD property in the setting of Riemannian geometry.
+Similarly, the MNHD property does not always hold for Riemannian manifolds [13, Appendix A].
+In the other direction, it was shown that the MNHD property holds on all flat tori [15] and some other Riemannian symmetric spaces including the $3$-dimensional hyperbolic space [13, Section 5 and Appendix B].
+It is still open whether the MNHD property holds on all Riemannian symmetric spaces of compact type and on all hyperbolic spaces.
+In a lattice in $\mathbb{R}^{n}$, the MNHD property reduces to a monotonicity property of the periodic Gaussian function.
+An extension of this monotonicity property to the non-spherical Gaussian case [16, Proposition 4.2] can be used to derive a monotonicity property of the normalized heat diffusion on a torus with respect to different flat Riemannian metrics [12, 16].
+Such an application also leads to a key ingredient in the framework of the so-called numerical cohomology theory [11, 14].
 2 Preliminaries
@@
 The graph Laplacian $L$ is a symmetric $n\times n$ matrix $L=(L(x,y))_{x,y\in V}$ defined by $L=\mathrm{Deg}-A$, where $\mathrm{Deg}$ is the diagonal vertex degree matrix and $A$ is the adjacency matrix.
-The constant vector $\mathbf{1}=(1,1,...,1)^{T}$ is an eigenvector of $L$ corresponding to the eigenvalue 00.
-Since $G$ is connected, the eigenvalue 00 is simple by the Perron-Frobenius theorem.
+The constant vector $\mathbf{1}=(1,1,...,1)^{T}$ is an eigenvector of $L$ corresponding to the eigenvalue 0.
+Since $G$ is connected, the eigenvalue 0 is simple by the Perron-Frobenius theorem.
 Let us denote by $\mathrm{sp}(L)$ the set of all distinct eigenvalues of $L$.
@@
 $L=\sum_{\lambda\in\mathrm{sp}(L)}\lambda P_{\lambda},$ (2)
-[…]
+where $P_{\lambda}$ is the projection operator onto the eigenspace associated with $\lambda$.
+The projection operator $P_{0}=\langle\cdot,\frac{\mathbf{1}}{\sqrt{n}}\rangle\frac{\mathbf{1}}{\sqrt{n}}$ corresponding to the simple eigenvalue 0 is an averaging operator, where we use $\langle\cdot,\cdot\rangle$ for the standard Euclidean inner product.
+In matrix form, we have $P_{0}=\frac{1}{n}J$, with $J$ being the $n\times n$ all-11 matrix.
 The heat kernel $H_{t}$ on the graph $G$ is defined in matrix form as $H_{t}:=e^{-tL}$. Accordingly, $H_{t}$ has the following spectral decomposition
@@
 2.3 Distance-regular graph with classical parameters
-A connected graph $G=(V,E)$ of diameter $D$ is distance-regular if and only if it is regular of degree $d$, and if there exist constants $c_{i},a_{i},b_{i}$, known as intersection numbers, such that for all $i=0,1,\dots,D$, and for any two vertices $x,y\in V$ at distance $i=d(x,y)$, among the neighbors of $y$, there are $c_{i}$ at distance $i-1$ from $x$, $a_{i}$ at distance $i$, and $b_{i}$ at distance $i+1$.
+A connected graph $G=(V,E)$ of diameter $D$ is distance-regular if and only if it is regular of degree $d$, and if there exist constants $c_{i},a_{i},b_{i}$, known as intersection numbers, such that for all $i=0,1,\dots,D$, and for any two vertices $x,y\in V$ at distance $i=d(x,y)$, among the neighbors of $y$, there are $c_{i}$ at distance i1i-1 from $x$, $a_{i}$ at distance $i$, and $b_{i}$ at distance $i+1$.
 It follows that $d=b_{0}$, and $a_{i}=d-b_{i}-c_{i}$ for all $i=0,1,\dots,D$.
@@
 To make $b_{i},c_{i}$ feasible parameters, there are necessary restrictions on the numbers $b,\alpha,\beta$.
-For example, the number $b$ must be an integer not equal to 00 or $-1$ [3, Proposition 6.2.1].
+For example, the number $b$ must be an integer not equal to 0 or 1-1 [3, Proposition 6.2.1].
 See also Lemma 4 below.
@@
 First, we derive from (11) and (12) the intersection numbers as follows.
-$\begin{array}[]{ll}b_{0}=d=(1+b+b^{2})\beta,&b_{1}=(b+b^{2})(\beta-\alpha),\\ b_{2}=b^{2}(\beta-\alpha(1+b)),&b_{3}=0,\\ c_{0}=0,&c_{1}=1,\\ c_{2}=(1+b)(1+\alpha),&c_{3}=(1+b+b^{2})(1+\alpha(1+b)),\\ a_{0}=0,&a_{1}=-1+b\alpha+b^{2}\alpha+\beta,\\ a_{2}=(1+b)(-1-\alpha+b^{2}\alpha+\beta),&a_{3}=(1+b+b^{2})(-1-\alpha-b\alpha+\beta).\\ \end{array}$ (13)
+[…]
 Then, we calculate the vertex number $n$ as below,
@@
 Lemma 5.
-For any three distinct numbers $\lambda_{i},\lambda_{j},\lambda_{k}\in\mathbb{R}$ with $\{i,j,k\}=\{1,2,3\}$ and any number $L(u,v),L^{2}(u,v)\in\mathbb{R}$, let $\Delta_{i}(u,v),i=1,2,3$ and $\Delta_{ij}(u,v)$ be defined in terms of $\lambda_{i},\lambda_{j},\lambda_{k},L(u,v),L^{2}(u,v)$ as in (7)-(10).
+For any three distinct numbers $\lambda_{i},\lambda_{j},\lambda_{k}\in\mathbb{R}$ with $\{i,j,k\}=\{1,2,3\}$ and any numbers $L(u,v),L^{2}(u,v)\in\mathbb{R}$, let $\Delta_{i}(u,v),i=1,2,3$ and $\Delta_{ij}(u,v)$ be defined in terms of $\lambda_{i},\lambda_{j},\lambda_{k},L(u,v),L^{2}(u,v)$ as in (7)-(10).
 Then it holds that
@@
 $\mathrm{III}:=$ $\lambda_{i}\lambda_{j}^{2}\lambda_{k}(\lambda_{j}-\lambda_{k})-\lambda_{i}^{2}\lambda_{j}\lambda_{k}(\lambda_{i}-\lambda_{k})+\lambda_{i}\lambda_{j}\lambda_{k}(\lambda_{i}+\lambda_{j}-\lambda_{k})(\lambda_{i}-\lambda_{j})$
-$=$ $0$
+$=$ $0.$
 Putting together yields that
@@
 Similarly as in Case $(i)$, we estimate
-$e^{-t(\lambda_{1}-\lambda_{3})}G^{\prime}(t)=\geq$ $\frac{\lambda_{1}\lambda_{2}}{n}\Delta_{1}(u,v)+\frac{\lambda_{2}\lambda_{1}}{n}\Delta_{2}(u,v)+\frac{\lambda_{3}(\lambda_{1}+\lambda_{2}-\lambda_{3})}{n}\Delta_{3}(u,v)$
+$e^{-t(\lambda_{1}-\lambda_{3})}G^{\prime}(t)\geq$ $\frac{\lambda_{1}\lambda_{2}}{n}\Delta_{1}(u,v)+\frac{\lambda_{2}\lambda_{1}}{n}\Delta_{2}(u,v)+\frac{\lambda_{3}(\lambda_{1}+\lambda_{2}-\lambda_{3})}{n}\Delta_{3}(u,v)$
 $+(\lambda_{3}-\lambda_{1})(\lambda_{2}-\lambda_{3})(\Delta_{13}(u,v)+\Delta_{23}(u,v))$
@@
 Remark 3.
-Observe from $L=dI-A$ that
+Observe from L=dIAL=dI-A that
 $L^{2}(x,y)$ $=d^{2}I(x,y)-2dA(x,y)+A^{2}(x,y)$
-=$=\begin{cases}d^{2}+d,&\text{if $x=y$,}\\ -2dA(x,y)+\sum_{w\neq x,y}A(x,w)A(w,y),&\text{if $x\neq y$.}\end{cases}$ (20)
+$=\begin{cases}d^{2}+d,&\text{if $x=y$,}\\ -2dA(x,y)+\sum_{w\neq x,y}A(x,w)A(w,y),&\text{if $x\neq y$.}\end{cases}$ (20)
 Therefore, for any non-adjacent vertices $u,v$, the condition $L^{2}(u,v)-L(u,v)(\lambda_{i}+\lambda_{j})\geq 0$ in Theorem 2 always holds.
@@
 Let $G=(V,E)$ be a distance-regular graph. Then we have for any $x,y\in V$ that
-$L^{2}\left(x,y\right)=\begin{cases}d^{2}+d,&\text{if}\;d\left(x,y\right)=0;\\ -2d+a_{1},&\text{if}\;d\left(x,y\right)=1;\\ c_{2},&\text{if}\;d\left(x,y\right)=2;\\ 0,&\text{if}\;d\left(x,y\right)\geq 3.\\ \end{cases}$ (21)
+[…]
 Proof.
@@
 $\gamma_{2}-\gamma_{1}$ $=b(\beta+1).$
-Recall that $b$ is an integer not equal to 00 or $-1$. Applying $\beta\geq 1+(1+b)\alpha\geq 1$ from Lemma 4, we deduce that
+Recall that $b$ is an integer not equal to 0 or 1-1. Applying $\beta\geq 1+(1+b)\alpha\geq 1$ from Lemma 4, we deduce that
 $\begin{cases}0<\gamma_{1}<\gamma_{2}<\gamma_{3},&\text{if $b\geq 1$;}\\ 0<\gamma_{2}<\gamma_{3}<\gamma_{1},&\text{if $b\leq-2$.}\end{cases}$ (23)
@@
 $=$ $(\beta-\alpha)(-db+b-\beta)<0.$
-Inserting into to (33), we obtain
+Inserting into (33), we obtain
 $\frac{n\Delta_{12}(u,v)}{C_{1}C_{2}(\lambda_{2}-\lambda_{1})}<0,$
@@
 $=$ $(db-d+\beta)(1+\alpha)-db(\beta-\alpha)\frac{b_{2}}{c_{3}}$
-$=$ $b\beta\left\{b^{2}-(1+b+b^{2})(\beta-\alpha)\frac{b_{2}}{c_{3}}.\right\}$
+$=$ $b\beta\left\{b^{2}-(1+b+b^{2})(\beta-\alpha)\frac{b_{2}}{c_{3}}\right\}.$
 Inserting the expressions $b_{2}=b^{2}(\beta-\alpha-b\alpha)$ and $c_{3}=(1+b+b^{2})(1+\alpha+b\alpha)$, we obtain
@@
 $\max_{1+(1+b)\alpha\leq\beta<1+(2+b)\alpha}g(\beta)\leq\max\{g(1+(1+b)\alpha),g(1+(2+b)\alpha)\}.$
-Since $-(1+b+b^{2})(1+\alpha)(1+\alpha+b\alpha)<0$, and the value $1+(1+b)\alpha$ lies in between $\alpha$ and $b+b\alpha+\alpha+1$, we have $g(1+(1+b)\alpha)<0$. Moreover, we check
+Since $(1+b+b^{2})(1+\alpha)(1+\alpha+b\alpha)<0$, and the value $1+(1+b)\alpha$ lies in between $\alpha$ and $b+b\alpha+\alpha+1$, we have $g(1+(1+b)\alpha)<0$. Moreover, we check
 $g(1+(2+b)\alpha)=(1+\alpha+b\alpha)(-b^{3}-(1+b+b^{2})-(1+b)(1+\alpha))<0.$
@@
 $d-\lambda_{1}=\theta_{2},\,\,d-\lambda_{2}=\theta_{3},\,\,\text{and}\,\,d-\lambda_{3}=\theta_{1}.$
-Using the calculations in the proof of Lemma 8, we check $(i)-(iii)$ as follows.
+Using the calculations in the proof of Lemma 8, we check (i)(iii)(i)-(iii) as follows.
 Case $(i)$: By (28), we have
@@
 It is natural to expect the monotonic normalized heat diffusion property holds for any distance regular graphs with classical parameters.
+Recall that Proposition 1 provides explicit expressions for all $D+1$ distinct Laplacian eigenvalues for any distance-regular graph with classical parameters $(D,b,\alpha,\beta)$.
+With this proposition in hand, the approach developed herein may pave the way for confirming the MNHD property for the cases where the number of distinct eigenvalues is greater than $4$.
 5 Further Discussion
@@
 Moreover, we have $L^{2}(u,v)-L(u,v)(\lambda_{2}+\lambda_{3})\geq 0$ for any two adjacent vertices $u$ and $v$.
-Recall that $L^{(}u,v)=-2d+a_{1}=-2d+(d-1-m\gamma)$ and $L(u,v)=-1$.
+Recall that $L(u,v)=-2d+a_{1}=-2d+(d-1-m\gamma)$ and L(u,v)=1L(u,v)=-1.
 We get $L^{2}(u,v)-L(u,v)(\lambda_{2}+\lambda_{3})=\frac{1}{2}(1+d+\gamma-m\gamma+M)\geq 0$, as desired.
@@
 $=$ $-\frac{m(d+1)}{4}\left\{(M+\gamma(1+m))^{2}-(d+1)^{2}-2\gamma(m+1)(1-d+\gamma(m+1)+M)\right\}$
-$=$ $0$
+$=$ $0.$
 Now we apply Theorem 2 to conclude that the function $r_{t}(u,v):=\frac{H_{t}(u,v)}{H_{t}(u,u)}$ is monotonically non-decreasing in $t$ for any two adjacent vertices $u$ and $v$.
@@
 and, hence, $\Delta_{13}(u,v)<0$. We apply Theorem 2 to conclude that the function $r_{t}(u,v)$ is monotonically non-decreasing in $t$ for any two vertices $u$ and $v$ with $d(u,v)=2.$
-Case $(iii)$: $d(u,v)=2$. Similar as in Case (i), we have
+Case $(iii)$: $d(u,v)=3$. Similar as in Case (i), we have
 $\Delta_{1}(u,v)/C_{1}=(\lambda_{2}-d)(\lambda_{3}-d)+d=\frac{1}{2}\left(1+d+\gamma+m\gamma+M\right)>0,$
@@
 Acknowledgement
+We are very grateful to the anonymous referee for suggestions that improved the readability of our manuscript.
 H.Z.
@@
 12431004.
+Availability of Data
+No data was used for the research described in the article.
+Declarations of Conflict of Interest
+The authors do not have any possible conflicts of interest.